(a) If the molar solubility of ScF3 at 25 oC is 6.81e-07 mol/L, what is the Ksp at this temperature? Ksp = (b) It is found that 0.00354 g of Ag2CO3 dissolves per 100 mL of aqueous solution at 25 oC. Calculate the solubility-product constant for Ag2CO3. Ksp = (c) The Ksp of ZnSe at 25 oC is 3.60e-26. What is the molar solubility of ZnSe? solubility =
Answer #1

Consider a reaction , ScF 3 (s)  php4fkmd2.png Sc 3+ (aq) + 3 F - (aq)

Equilibrium constant for above reaction is K sp = [Sc 3+ ] [F - ] 3

If S is the solubility of ScF 3  in mol / L , then [Sc 3+ ] = S mol / L and [F - ] = 3 S mol / L.

phpkakz1j.png K sp = [Sc 3+ ] [F - ] 3 = S phpIpVINn.png (3 S ) 3

K sp = S phpIpVINn.png 27 S 3

K sp = 27 S 4

We have, S = 6.81 phpIpVINn.png 10 -07 M

phpkakz1j.png K sp = 27 (6.81 phpIpVINn.png 10 -07) 4

K sp =5.81 phpIpVINn.png 10 -24

b) Solubility of Ag2CO3 = 0.00354 g / 100 ml

We have relation , 1 L = 10 ( 100 ml )

phpkakz1j.png Solubility of Ag2CO3 = 0.00354 g / 100 ml  phpIpVINn.png ( 10 ( 100 ml ) / 1 L ) = 0.0354 g / L

Molar Mass of Ag2CO3 = ( 2 phpIpVINn.png 107.87) + 12.01 + ( 3 phpIpVINn.png 16.00) = 275.75 g / mol

Molar solubility of Ag2CO3 = ( 0.0354 g / L) / ( 275.75 g / mol) = 1.284 phpIpVINn.png 10 -04 mol / L

Consider a reaction , Ag2CO 3 (s)  phpXcbzcB.png 2 Ag + (aq) + CO 32- (aq)

Equilibrium constant for above reaction is K sp = [Ag + ] 2 [CO 32- ]

If S is the solubility of Ag2CO 3 in mol / L , then [Ag + ] = 2 S mol / L and [CO 32- ] = S mol / L.

phpOLHCnH.png K sp =[Ag + ] 2 [CO 32- ]= (2 S ) 2phpJC7pwu.png S = 4 S 2 phpJC7pwu.png S = 4 S 3

We have calculated S = 1.284 phpIpVINn.png 10 -04 mol / L.

phpOLHCnH.png K sp = 4 (1.284 phpIpVINn.png 10 -04 )  3 = 8.467 phpIpVINn.png 10 -12

ANSWER : K sp of Ag2CO 3 = 8.467 phpIpVINn.png 10 -12

C)

Consider a reaction , ZnSe (s)  phpzqEra9.png Zn 2+ (aq) + Se 2- (aq)

Equilibrium constant for above reaction is K sp = [Zn 2+ ] [Se 2- ] = 3.60 phpnvTozz.png 10 -26

If S is the solubility of ZnSe in mol / L , then [Zn 2+ ] = S mol / L and [Se 2- ] = S mol / L.

php0JLI9b.png K sp = [Zn 2+ ] [Se 2- ] = S phpAVDvU0.png S = S 2

S = K sp

S = V3.60 x 10e - 26

S = 1.897 phpAVDvU0.png 10 -13 M

ANSWER : Molar solubility of ZnSe is 1.897 phpAVDvU0.png 10 -13 M